OpenStudy (anonymous):

Can you guys help me with this question? x 4 - = - x+2 5 Thank you please answer fast!

OpenStudy (nottim):

Try using the Equation button, or draw it out. Looks confusing.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

^ ? OKAYY

OpenStudy (nottim):

http://www.youtube.com/watch?v=iZst_2xJHAI An appropriate Theme Song for the Bunny Person.

OpenStudy (katrinakaif):

Cross mulitply 5x = 4(x+2) Distribute 4 5x = 4x + 8 Bring all the x's to one side (by substracting 4x from 5x) 5x - 4x = 8 1x = 8 Divide by 1 x = 8

OpenStudy (karatechopper):

jeepers ur fast!

OpenStudy (nottim):

Agreed. too fast. Do you KNOW OF THE MATRIX, MA'AM?

OpenStudy (anonymous):